### v2

# Unified Molecule-Text Language Model with Discrete Token Representation

Abstract
The remarkable success of Large Language Models (LLMs) across diverse tasks has driven the research community to extend their capabilities to molecular applications. However, most molecular LLMs employ adapter-based architectures that fail to equally integrate molecule and text modalities and lack explicit supervision signals for the molecular modality. To address these issues, we introduce UniMoT, a Unified Molecule-Text LLM adopting a tokenizer-based architecture that expands the vocabulary of LLMs with molecule tokens. Specifically, we introduce a Vector Quantization-driven tokenizer that incorporates a Q-Former to bridge the modality gap between molecule and text. This tokenizer transforms molecular structures into sequences of tokens exhibiting causal dependency, thereby encapsulating both high-level molecular features and textual information. Equipped with this tokenizer, UniMoT unifies molecule and text modalities under a shared token representation and an autoregressive training paradigm. This enables the model to process molecular structures as a distinct linguistic system and generate them in textual form. Through a four-stage training scheme, UniMoT functions as a multi-modal generalist capable of performing both molecule-to-text and text-to-molecule tasks. Extensive experiments demonstrate that UniMoT achieves state-of-the-art performance across a wide range of molecule comprehension and generation tasks.
1 Introduction
The incredible capabilities of Large Language Models (LLMs) Brown et al. (2020); Touvron et al. (2023) have led to their widespread use as versatile tools for completing diverse real-world tasks. This success has sparked interest in Multi-modal LLMs Zhan et al. (2024), which aim to enhance LLMs by enabling them to process multi-modal inputs and outputs. In fields like molecular science, Multi-modal LLMs present new opportunities by seamlessly integrating molecular data with textual information, opening up fresh possibilities for more efficient and accurate research and development.Prior research efforts Liang et al. (2023); Fang et al. (2023); Cao et al. (2023); Liu et al. (2023b); Li et al. (2024) have focused on adapting LLMs to molecular tasks, resulting in the development of molecular LLMs. These molecular LLMs can analyze molecule structures Liu et al. (2023b); Cao et al. (2023), address drug-related inquiries Liang et al. (2023), assist in synthesis and retrosynthesis planning Fang et al. (2023), support drug design Fang et al. (2023), and more.
(a) Projection-Based Architecture.
(b) Q-Former-Based Architecture.
(c) Tokenizer-Based Architecture.
Figure 1: Comparisons among different molecular LLMs. 1(a) and 1(b) are adapter-based architectures that do not treat molecule and text modalities equally and lack a supervision signal for the molecule modality. 1(c) is our proposed tokenizer-based architecture, where molecules are presented in the same discrete token representation as that of text.
Prevalent molecular LLMs often use adapter-based architectures, such as linear projection Liang et al. (2023); Cao et al. (2023) or Q-Former Liu et al. (2023b); Li et al. (2024), to map molecule features into the LLM’s semantic space (Figure 1(a), Figure 1(b)). While effective in molecular comprehension and molecule-to-text generation, these models struggle with text-to-molecule generation. This is due to the reliance on adapters that require LLMs to directly generate SMILES strings Weininger (1988), a text-based representation of molecular structures. These architectures depend on strong alignment between SMILES and text, but as shown in Figure 1(a) and Figure 1(b), molecule and text modalities are not treated equally, with insufficient supervision for the molecular side, making alignment difficult.
Discretizing continuous molecule features into discrete molecule tokens offers a promising solution for conducting both molecule-to-text and text-to-molecule generation tasks. By treating tokens from different modalities equally, we can predict the next molecule or text token in an autoregressive manner. However, directly discretizing molecule features poses several challenges: (i) This approach results in long sequences, with lengths equivalent to the number of atoms in a batch; (ii) Molecule tokens derived from molecule features lack left-to-right causal dependency, which conflicts with the unidirectional attention mechanism in LLMs; (iii) Molecule features lack textual information, hindering effective molecule-text interactions and alignment.
To this end, we present UniMoT, a Unified Molecule-Text LLM that adopts a tokenizer-based architecture, integrating molecule comprehension and generation, as depicted in Figure 1(c). A pivotal aspect of UniMoT’s architecture is the molecule tokenizer for transforming molecules into molecule tokens. We introduce a Vector Quantization-driven Van Den Oord et al. (2017) tokenizer, which incorporates a Q-Former Li et al. (2023) to bridge the modality gap between molecules and text. Specifically, we incorporate causal masks for the queries, enabling the Q-Former to generate a causal sequence of queries compatible with the unidirectional attention in LLMs. The sequence of queries is subsequently quantized into a sequence of molecule tokens using a learnable codebook. The molecule tokens encapsulate high-level molecular and textual information, which are then aligned with the latent space of a pretrained generative model via an MLP adapter.
Pretrained LLMs can integrate the molecule tokenizer by treating molecule tokens as new words and constructing a molecule vocabulary through mapping the learned codebook. We adopt the unified discrete token representation for molecules and text, coupled with the unified next-token-prediction training paradigm of LLM. This unification of representation and training paradigm enables effective molecule-text interactions and alignment through molecule-to-text and text-to-molecule autoregressive pretraining. For molecule generation tasks, UniMoT generates molecule tokens in an autoregressive manner rather than producing SMILES strings, and these molecule tokens can then be decoded into molecules using the generative model.
Our contributions can be summarized as follows:
  • We introduce a molecule tokenizer specifically designed for LLMs, enabling the tokenization of molecules into short sequences of tokens with causal dependency. These tokens encapsulate high-level molecular and textual information and can be decoded into desired molecules during inference.
  • We present UniMoT, a unified molecule-text LLM that adopts a tokenizer-based architecture instead of traditional adapter-based architectures. UniMoT unifies the modalities of molecule and text under a shared token representation and an autoregressive training paradigm. Following a four-stage training scheme, UniMoT effectively achieves molecule-text alignment.
  • UniMoT exhibits remarkable capabilities in multi-modal comprehension and generation. Extensive experiments show that UniMoT achieves state-of-the-art performance across a wide range of comprehension and generation tasks, while also offering a new perspective on molecule generation.
2 Related Works
Multi-modal Large Language Models.
Multi-modal Large Language Models (LLMs): Current multi-modal LLMs are typically built on a pre-trained LLM backbone and can understand multiple modalities. LLaVA Liu et al. (2024a) connects the image encoder to the LLM using a simple linear projection, while BLIP-2 Li et al. (2023) extracts high-level features from images with CLIP Radford et al. (2021) and uses Q-Former to reduce image token counts. While these models excel at multi-modal comprehension, they often lack focus on multi-modal generation. To address this, recent work unifies multi-modal comprehension and generation, such as SEED-LLaMA Ge et al. (2023) and AnyGPT Zhan et al. (2024), which unify processing across different modalities. Inspired by these advances, we introduce a tokenizer-based architecture in the molecule-text domain, converting molecular features into tokens compatible with LLMs.
Molecular Large Language Models.
The recent emergence of Vision Large Language Models (VLLMs) Li et al. (2023) has catalyzed advancements in molecular LLMs, which encompass both single modality and multi-modality approaches. In the single modality domain, researchers are exploring diverse molecule representations, such as 1D sequences like SMILES strings Irwin et al. (2022), 2D molecule graphs You et al. (2020), 3D geometric conformations You et al. (2020), and textual information from the literature Taylor et al. (2022). In the multiple modalities domain, various innovative approaches are being employed. MolT5 Edwards et al. (2022), a T5-based Raffel et al. (2020) model, is designed for SMILES-to-text and text-to-SMILES translations. Other works, such as MoMu Su et al. (2022), MoleculeSTM Liu et al. (2023a), and GIT-Mol Liu et al. (2024b), leverage cross-modal contrastive learning to align the representation spaces of molecules and text. Additionally, some studies Cao et al. (2023); Liang et al. (2023); Liu et al. (2023b); Li et al. (2024) use multi-modal learning architectures to develop molecular LLMs, which often adopt adapter-based architectures. However, these methods do not treat molecule and text modalities equally and lack a supervision signal for the molecule modality, limiting model capacity and effectiveness.
Figure 2: Illustration of our proposed molecule tokenizer. The tokenizer generates discrete molecule tokens, which can be fed into LLMs for downstream tasks. The generated molecule tokens can be decoded into molecules using the adapter and the SMILES decoder during inference.
Vector Quantization.
Vector Quantization (VQ) Gray (1984) is a widely used technique in generative models. VQ-VAE Van Den Oord et al. (2017) converts an image into a set of discrete codes within a learnable discrete latent space by learning to reconstruct the original image. VQ-GAN Yu et al. (2021) enhances the generation quality by leveraging adversarial and perceptual objectives. In the context of molecules, VQ has been effectively applied to quantize molecule features. For example, DGAE Boget et al. (2023) introduces a VQ model specifically for molecules, where molecules are encoded into discrete latent codes. Mole-BERT Xia et al. (2022) uses VQ to rethink the pre-training of GNNs for molecular tasks. IMoLD Zhuang et al. (2024) proposes using VQ to enhance invariant molecule representations, and VQSynergy Wu et al. (2024) demonstrates the use of VQ for drug discovery.
3 Method
Our objective is to leverage the reasoning and generation capabilities of LLMs to enhance the comprehension and generation of molecule and text data. To achieve this, we focus on representing these modalities uniformly within the token representation, utilizing the next-token-prediction training paradigm of LLMs. As illustrated in Figure 2, we introduce a molecule tokenizer (Section 3.1) designed to transform molecules into molecule tokens by learning to reconstruct the input molecule. The molecule sequence can then be concatenated with the text sequence to form a multi-modal sequence, which is fed into an LLM for molecule-to-text and text-to-molecule autoregressive pretraining (Section 3.2), as illustrated in Figure 3. The LLM vocabulary is expanded with molecule tokens mapped from the learned codebook. We introduce a four-stage training scheme for UniMoT (Section 3.3) comprising Causal Q-Former pretraining, molecule tokenizer pretraining, unified molecule-text pretraining, and task-specific instruction tuning. UniMoT is capable of performing molecule comprehension and generation tasks following the training scheme.
(a) Molecule-to-Text Autoregression.
(b) Text-to-Molecule Autoregression.
Figure 3: Illustration of the multi-modal autoregressive pretraining on molecule-text datasets. UniMoT excels in multi-modal comprehension and generation tasks, enabled by the unified LM objective. $T$ represents the size of the text vocabulary.
3.1 Molecule Tokenizer for LLMs
Molecule Encoder.
We represent the structural information of a molecule as a graph, denoted by $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of atoms and $|\mathcal{V}|=N$ is the number of atoms. The task of the molecule encoder is to extract molecule features that are context-aware and encompass diverse local neighborhood structural information. By employing a molecule encoder, we obtain molecule features $\mathbf{X}\in\mathbb{R}^{N\times F}$, where $F$ denotes the dimensionality of the feature vector for each atom.
Causal Q-Former.
We employ a Q-Former model introduced by BLIP-2 Li et al. (2023) to generate queries $\mathbf{Z}=\{\boldsymbol{z}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times d}$ containing high-level molecular and textual information, where $M$ represents the number of queries and $d$ denotes the dimension of queries. The Q-Former operates as a query-based transformer that utilizes learnable queries $\{\boldsymbol{z}_{i}\}_{i=1}^{M}$ to interact with molecule features $\mathbf{X}$ extracted by the molecule encoder. Specifically, we incorporate causal masks into the queries, ensuring that they only interact with preceding queries. This ensures the sequence of queries maintains a causal dependency, aligning with the unidirectional requirements of LLMs operating on text sequence. Details regarding the Causal Q-Former can be found in Appendix A.
Vector Quantization.
The Causal Q-Former converts molecules and text into a causal sequence of queries. Subsequently, the causal sequence of queries $\{\boldsymbol{z}_{i}\}_{i=1}^{M}$ is quantized into a causal sequence of molecule tokens $\{s_{i}\}_{i=1}^{M}$ by identifying the closest neighbor in a learnable codebook $\mathcal{C}=\{\boldsymbol{c}_{i}\}_{i=1}^{K}$, where $K$ represents the size of the codebook. The codebook is randomly initialized and optimized during pretraining. Specifically, token $s_{i}$ is determined as follows:
$s_{i}=\operatorname{argmin}_{j\in\{1,\cdots,K\}}\left\|\boldsymbol{z}_{i}-\boldsymbol{c}_{j}\right\|_{2},\quad\text{for}\quad i=1,2,\cdots,M.$
Intuitively, the query $\boldsymbol{z}_{i}$ is quantized to the closest neighbor $\boldsymbol{c}_{s_{i}}$ in the codebook. As the vector quantization process is non-differentiable, we adopt the straight-through estimator Bengio et al. (2013) to train the Causal Q-Former by copying the gradient from the molecule tokens to the queries, as shown in Figure 2. The resulting embeddings of molecule tokens $\{s_{i}\}_{i=1}^{M}$, denoted as $\mathbf{C}=\{\boldsymbol{c}_{s_{i}}\}_{i=1}^{M}$, are subsequently utilized for reconstructing molecules.
Molecule Reconstruction.
An MLP adapter $\psi$ needs to be trained to align the discrete latent space of molecule tokens with the continuous latent space of a molecular generative model for molecule reconstruction. This can be represented as $\mathbf{X}_{R}=\psi(\mathbf{C})$, where $\mathbf{X}_{R}$ denotes the embeddings for reconstruction. To achieve alignment, we minimize the Mean Squared Error (MSE) loss between $\mathbf{X}_{R}$ and the SMILES Weininger (1988) embeddings $\mathbf{X}_{S}$ produced by the pretrained SMILES encoder. Subsequently, we can reconstruct the molecule from $\mathbf{X}_{R}$ using the pretrained SMILES decoder. The training loss of the tokenizer is expressed as follows:
$\mathcal{L}_{\text{Tokenizer}}=\left\|\mathbf{X}_{R}-\mathbf{X}_{S}\right\|_{2}^{2}+\frac{1}{M}\sum_{i=1}^{M}\left\|\operatorname{sg}\left[\boldsymbol{z}_{i}\right]-\boldsymbol{c}_{s_{i}}\right\|_{2}^{2}\\ +\frac{\beta}{M}\sum_{i=1}^{M}\left\|\operatorname{sg}\left[\boldsymbol{c}_{s_{i}}\right]-\boldsymbol{z}_{i}\right\|_{2}^{2}.$ (2)
Here, the first term represents the alignment loss, the second term is a codebook loss aimed at updating the codebook embeddings, and the third term is a commitment loss that encourages the query to stay close to the chosen codebook embedding. $\operatorname{sg}[\cdot]$ denotes the stop-gradient operator, and the hyperparameter $\beta$ is set to 0.25.
3.2 Unified Molecule-Text Language Model
Expanding Vocabulary.
Employing the molecule tokenizer, a molecule can be tokenized into a molecule sequence $\{s_{i}\}_{i=1}^{M}$ with causal dependency. The molecule sequence can be concatenated with the text sequence to form a multi-modal sequence $\{u_{i}\}_{i=1}^{L}$, where $L$ is the length of the multi-modal sequence. To facilitate the representation of the multi-modal sequence, we construct the molecule vocabulary $\mathcal{V}^{m}=\{\boldsymbol{v}^{m}_{i}\}_{i=1}^{K}$, which maintains the order of the molecule codebook $\mathcal{C}=\{\boldsymbol{c}_{i}\}_{i=1}^{K}$. Additionally, $\mathcal{V}^{m}$ includes several special tokens such as boundary indicators, e.g., [MOL] and [/MOL], to mark the beginning and end of the molecule sequence. Next, we merge the original text vocabulary $\mathcal{V}^{t}=\{\boldsymbol{v}^{t}_{i}\}_{i=1}^{T}$ with the molecule vocabulary $\mathcal{V}^{m}$. The unified molecule-text vocabulary $\mathcal{V}=\{\mathcal{V}^{m},\mathcal{V}^{t}\}$ facilitates joint learning from molecules and text under a unified next-token-prediction objective. As the vocabulary is expanded, the corresponding embeddings and prediction layers also need to be extended, with the newly introduced parameters initialized randomly.
[…]
KV-PLM 70.50 72.12 55.03 59.83 89.17 54.63 65.40 78.50
AttrMask 67.79 75.00 63.57 58.05 75.44 73.76 75.44 80.28
InfoGraph 64.84 76.24 62.68 59.15 76.51 72.97 70.20 77.64
MolCLR 67.79 75.55 64.58 58.66 84.22 72.76 75.88 71.14
GraphMVP 68.11 77.06 65.11 60.64 84.46 74.38 77.74 80.48
MoleculeSTM 69.98 76.91 65.05 60.96 92.53 73.40 76.93 80.77
InstructMol (Vicuna-7B) 70.00 74.67 64.29 57.80 91.48 74.62 68.90 82.30
UniMoT (Llama-2-7B) 71.37 76.43 65.78 59.79 92.89 75.97 78.49 83.69
Table 1: ROC-AUC (%) of molecular property prediction task (classification) on the MoleculeNet Wu et al. (2018) datasets. Bold indicates the best performance and underline indicates the second best performance.
Unified Molecule-text Modeling.
The multi-modal sequence $\{u_{i}\}_{i=1}^{L}$ is fed into the pretrained LLM for performing multi-modal autoregression. UniMoT adopts the general Language Modeling (LM) objective to directly maximize the log-likelihood of the data distribution:
$\mathcal{L}_{\text{LM}}=-\sum_{u\in\mathcal{D}}\sum_{i\in\mathcal{I}}\log p\left(u_{i}\mid u_{1},\cdots,u_{i-1};\theta\right),$ (3)
where $\mathcal{D}$ represents the dataset, $\mathcal{I}$ represents the set of indices of the generation target, and $\theta$ denotes the parameters of the LLM. The unification of representation and training paradigm for molecules and text enhances the abilities of LLMs to understand molecule-text interactions and alignment. UniMoT can interpret molecules similar to understanding a foreign language, and generate them as if they were text. We conduct autoregressive pretraining on molecule-to-text and text-to-molecule tasks to enhance the molecule comprehension and generation capabilities.
Molecule-to-Text Autoregression.
While structural information is embedded in molecule features and captured by the molecule tokens through the tokenizer, we also aim to incorporate sequential information of molecules for better comprehension. Therefore, we concatenate the molecule sequence $\{s_{i}\}_{i=1}^{M}$ with the SMILES Weininger (1988) sequence and a prompt to form the multi-modal input sequence $\{u_{i}\}_{i=1}^{L}$, as illustrated in Figure 3(a). The corresponding molecule caption is used as the generation target.
Text-to-Molecule Autoregression.
For molecule generation, a prompt and the molecule caption are concatenated, with a [MOL] token appended to signify the beginning of the molecule sequence, as illustrated in Figure 3(b). The molecule sequence $\{s_{i}\}_{i=1}^{M}$ produced by the tokenizer is used as the generation target. During inference, given a prompt and the molecule caption, the output molecule sequence can be decoded into the desired molecule by the pretrained adapter and SMILES decoder.
3.3 Training Strategy
The training strategy for UniMoT is structured across four stages. Stage-1 focuses on Causal Q-Former pretraining with tailored objectives. In Stage-2, the molecule tokenizer is optimized using the frozen encoders and decoder. Stage-3 integrates the tokenizer with a language model for multi-modal comprehension and generation. Finally, Stage-4 fine-tunes UniMoT for specific tasks, aligning it with human instructions and optimizing performance for various molecular applications. More details regarding the training process can be found in Appendix C.
Stage-1: Causal Q-Former Pretraining.
We connect the molecule encoder and Causal Q-Former, leveraging the pretrained MoleculeSTM molecule encoder Liu et al. (2023a). The molecule encoder remains frozen while only the Causal Q-Former is updated. Both queries and text inputs are used, while only queries serve as input in subsequent stages. In our experiments, we utilize 16 queries. We employ three tailored objectives for the pretraining of the Causal Q-Former: Molecule-Text Contrastive Learning (MTC), Molecule-Text Matching (MTM), and Molecule-grounded Text Generation (MTG). The details of these objectives can be found in Appendix A.
Stage-2: Molecule Tokenizer Pretraining.
We connect the Causal Q-Former with subsequent blocks and use the objective defined in Equation (2). We employ the pretrained ChemFormer Irwin et al. (2022) as the generative model. Specifically, we leverage the SMILES encoder and the SMILES decoder provided by ChemFormer. The molecule codebook size is set to $K=2048$. As shown in Figure 2, we keep the molecule encoder, the SMILES encoder, and the SMILES decoder frozen, while updating the Causal Q-Former, the learnable codebook, and the adapter.
Stage-3: Unified Molecule-Text Pretraining.
We integrate the molecule tokenizer with the LLM using the unified vocabulary of molecule tokens and text tokens. We employ the LM objective defined in Equation (3) to pretrain the LLM. Pretraining involves molecule-to-text autoregression and text-to-molecule autoregression, aimed at enhancing UniMoT’s multi-modal comprehension and generation capabilities. To enhance efficiency, we train the LLM using low-rank adaptation (LoRA) Hu et al. (2021).
Stage-4: Task-Specific Instruction Tuning.
UniMoT is fine-tuned on seven comprehension and generation tasks: molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. We also utilize LoRA to improve efficiency. This stage ensures UniMoT can accurately interpret and respond to human instructions, making it versatile and effective for diverse molecular tasks.
4 Experiments
4.1 Molecule Comprehension Tasks
Molecular Property Prediction Task.
The goal of molecular property prediction is to forecast a molecule’s intrinsic physical and chemical properties. For the classification task, we incorporate eight binary classification datasets from MoleculeNet Wu et al. (2018). Models are tasked with generating a single prediction (“yes” or “no”). We compare UniMoT with the following baselines: KV-PLM Zeng et al. (2022), AttrMask Hu et al. (2019), InfoGraph Sun et al. (2019), MolCLR Wang et al. (2021), GraphMVP Liu et al. (2019), MoleculeSTM Liu et al. (2023a), and InstructMol Cao et al. (2023). The ROC-AUC (%) results on the MoleculeNet datasets are shown in Table 1. The performance of the regression task of molecular property prediction is provided in Appendix D. Compared to traditional graph learning methods and molecular LLMs like InstructMol Cao et al. (2023), UniMoT demonstrates consistent improvements across the eight datasets, indicating its robust molecule comprehension abilities.
[…]
MolT5-Small (T5-Small) 22.5 15.2 30.4 13.5 20.3 24.0
MolT5-Base (T5-Base) 24.5 16.6 32.2 14.0 21.4 26.1
MolT5-Large (T5-Large) 25.9 17.3 34.1 16.4 23.4 28.0
MoMu-Small (T5-Small) 22.9 16.0 31.0 13.7 20.8 24.4
MoMu-Base (T5-Base) 24.7 16.8 32.5 14.6 22.1 27.2
MoMu-Large (T5-Large) 26.3 18.0 34.8 16.9 24.8 28.7
InstructMol (Vicuna-7B) 18.9 11.7 27.3 11.8 17.8 21.3
MolCA (OPT-125M) 25.9 17.5 34.4 16.6 23.9 28.5
MolCA (OPT-1.3B) 28.6 21.3 36.2 21.4 29.7 32.6
3D-MoLM (Llama-2-7B) 30.3 22.5 36.8 22.3 31.2 33.1
UniMoT (Llama-2-7B) 31.3 23.8 37.5 23.7 33.6 34.8
Table 2: Performance (%) of molecule captioning task on the PubChem Kim et al. (2023) dataset. Bold indicates the best performance and underline indicates the second best performance.
Molecule Captioning Task.
The molecule captioning task involves generating a comprehensive description of a molecule. We compare UniMoT with several baselines: MolT5 Edwards et al. (2022), MoMu Su et al. (2022), InstructMol Cao et al. (2023), MolCA Liu et al. (2023b), and 3D-MoLM Li et al. (2024). BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005) are adopted as evaluation metrics. UniMoT is evaluated for molecule captioning on the PubChem Kim et al. (2023) and ChEBI-20 Edwards et al. (2022) datasets. Performance on the PubChem dataset is shown in Table 2, while the performance on the ChEBI-20 dataset and some concrete examples are presented in Appendix D. The ChEBI-20 dataset replaces molecular names with “the molecule” to focus on properties. However, predicting molecular names reflects the model’s structural understanding, so we conducted the main experiments on PubChem.
From Table 2, we observe that UniMoT consistently outperforms the baselines by a significant margin on the PubChem Kim et al. (2023) dataset. This task is more complex than classification or regression, providing a robust measure of the model’s molecule comprehension abilities. Notably, our proposed tokenizer-based architecture surpasses the projection-based architecture (such as InstructMol Cao et al. (2023)), Q-Former-based architecture (such as MolCA Liu et al. (2023b) and 3D-MoLM Li et al. (2024)), and models trained with contrastive learning strategies (such as MoMu Su et al. (2022)). This demonstrates that the tokenizer-based architecture achieves better molecule-text alignment through autoregressive molecule-to-text and text-to-molecule pretraining compared to other architectures. Details and More Results of Experiments can be found in Appendix D.
4.2 Molecule Generation Tasks
We employ molecule generation tasks, which encompass caption-guided molecule generation Fang et al. (2023), reagent prediction Fang et al. (2023), forward reaction prediction Fang et al. (2023), and retrosynthesis Fang et al. (2023). Caption-guided molecule generation involves generating molecular structures based on textual descriptions. Reagent prediction entails determining suitable reagents given reactants and products. Forward reaction prediction involves predicting probable products given specific reactants and reagents. Retrosynthesis involves deconstructing a target molecule into simpler starting materials. We compare UniMoT with the following baselines: Llama Touvron et al. (2023), Vicuna Chiang et al. (2023), Mol-Instructions Fang et al. (2023), and InstructMol Cao et al. (2023). The metrics used to evaluate molecule generation tasks include Exact Match, BLEU Papineni et al. (2002), Levenshtein Distance Levenshtein and others (1966), RDKit Fingerprint Similarity Landrum and others (2006), MACCS Fingerprint Similarity Durant et al. (2002), and Morgan Fingerprint Similarity Morgan (1965). These metrics evaluate structural similarity between generated and target molecules, along with Validity Kusner et al. (2017), which assesses the proportion of chemically valid molecules generated.
We utilize the Mol-Instructions Fang et al. (2023) benchmark to evaluate the generation capabilities of UniMoT. The results of caption-guided molecule generation and reagent prediction are presented in Table 3, and the results of other tasks are in Appendix D. The caption-guided molecule generation task, the reverse of molecule captioning, is conducted using the PubChem Kim et al. (2023) dataset, while the other tasks utilize the USPTO Fang et al. (2023) dataset. As the baselines generate SMILES strings and then convert them to molecules, UniMoT directly leverages the generated molecule tokens and obtains their embeddings from the learned codebook. These embeddings can be decoded to desired molecules through the pretrained adapter and SMILES decoder. As shown in Table 3, UniMoT generates valid molecules with a higher degree of similarity to the target molecules compared to the baselines. This is because UniMoT can generate molecules as if they were text, which is fundamentally different from adapter-based architectures. UniMoT demonstrates strong generation capabilities and offers a new perspective on these tasks.
[…]
Caption-guided Molecule Generation
Llama 0.000 0.003 59.864 0.005 0.000 0.000 0.003
Vicuna 0.000 0.006 60.356 0.006 0.001 0.000 0.001
Mol-Instructions 0.002 0.345 41.367 0.231 0.412 0.147 1.000
MolT5 0.112 0.546 38.276 0.400 0.538 0.295 0.773
UniMoT 0.237 0.698 27.782 0.543 0.651 0.411 1.000
Reagent Prediction
Llama 0.000 0.003 28.040 0.037 0.001 0.001 0.001
Vicuna 0.000 0.010 27.948 0.038 0.002 0.001 0.007
Mol-Instructions 0.044 0.224 23.167 0.237 0.364 0.213 1.000
InstructMol 0.129 0.610 19.664 0.444 0.539 0.400 1.000
UniMoT 0.167 0.728 14.588 0.549 0.621 0.507 1.000
Table 3: Performance of molecule generation tasks on the Mol-Instructions Fang et al. (2023) benchmark, including caption-guided molecule generation and reagent prediction. Bold indicates the best performance, and underline indicates the second best performance.
4.3 Ablation Studies
Cross-Modal Projector.
Projector Input to LLM BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
Projection Layer Molecule Emb. 19.3 12.1 27.9 12.3 18.1 21.5
Q-Former Query Emb. 28.6 21.3 36.2 21.4 29.7 32.6
Causal Q-Former Causal Emb. 32.8 25.2 39.2 24.8 35.3 36.5
Causal Q-Former Causal Tokens 31.3 23.8 37.5 23.7 33.6 34.8
(a) Ablation study on the projector and representation form.
Architecture Codebook Size BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
Llama-2-7B 512 28.7 20.5 33.2 20.7 29.6 30.2
Llama-2-7B 1024 29.5 21.3 34.5 21.8 30.9 31.5
Llama-2-7B 2048 31.3 23.8 37.5 23.7 33.6 34.8
Llama-2-7B 4096 31.1 23.6 37.1 23.5 33.2 34.3
(b) Ablation study on the codebook size.
Table 4: Ablation studies on the molecule captioning task using the PubChem dataset.
We conducted an ablation study on the cross-modal projector, with the results on the molecule captioning task shown in Table 4(a). The linear projection demonstrated the worst performance, indicating that the molecule features lack textual information, thus hindering effective molecule-text interactions and alignment. Additionally, we compared the performance of a Q-Former with bidirectional self-attention to a Causal Q-Former with causal self-attention in the second and third rows. The results show that queries with causal dependency outperform those with bidirectional dependency. This demonstrates that input with left-to-right causal dependency aligns with the unidirectional attention mechanism in LLMs, leading to improved performance.
Discrete vs. Continuous Representation.
We compared the performance of continuous causal embeddings and discrete tokens, quantized from causal embeddings, as inputs to LLMs in the third and fourth rows of Table 4(a). Continuous embeddings demonstrate better performance than discrete tokens in understanding molecules. This result is reasonable since the quantization process causes information loss in discrete tokens. However, we still use discrete token representation to facilitate the autoregressive training paradigm of LLMs, which supports the unification of comprehension and generation tasks. To achieve this unification, we unavoidably sacrifice some performance in comprehension tasks.
Codebook Size.
We conducted experiments with different molecule codebook sizes and reported the performance on the molecule captioning task. The performance is shown in Table 4(b). The results demonstrate that the codebook size of 2048 consistently provides the best performance. This choice balances model complexity and performance. A larger codebook could capture more subtle interactions between molecules and text. However, there may be some codes that are not often used. A smaller codebook may result in nearby embeddings being assigned the same code, which reduces the granularity of the representation. More ablation studies are presented in Appendix E.
5 Conclusion
This work introduces UniMoT, a framework that unifies the modalities of molecules and text. By adopting a tokenizer-based architecture, UniMoT addresses previous limitations where the molecule and text modalities are not treated equally. The molecule tokenizer converts molecules into sequences of discrete tokens, embedding high-level molecular and textual information. The LLM vocabulary is expanded with molecule tokens mapped from a learned codebook. Through a four-stage training scheme, UniMoT has become a versatile multi-modal LLM, capable of handling both molecule-to-text and text-to-molecule tasks. Extensive empirical evaluations show that UniMoT achieves state-of-the-art performance across diverse molecule comprehension and generation tasks.
Although UniMoT excels in molecule-to-text and text-to-molecule tasks, it has yet to be extensively tested on more complex tasks like molecule editing, which require precise structural modifications. Additionally, limited annotated data in the molecular domain restricts UniMoT’s training, hindering its ability to fully learn and generalize molecular structures and properties. To improve its effectiveness, addressing data scarcity is crucial. Furthermore, expanding evaluations to include a broader range of real-world scenarios will offer a more comprehensive understanding of the model’s robustness and generalizability.
Acknowledgments
Q. Yao is supported by National Natural Science Foundation of China (under Grant No.92270106), Beijing Natural Science Foundation (under Grant No.4242039), and CCF-Zhipu Large Model Innovation Fund (CCF-Zhipu202402)
Appendix A Details of Causal Q-Former
The Q-Former operates as a query-based transformer that utilizes learnable query vectors to interact with molecule features extracted by a frozen encoder. These queries are essential for extracting relevant information from the molecule features. The Q-Former comprises both a molecule transformer and a text transformer, sharing self-attention layers. The molecule transformer incorporates cross-attention layers between self-attention and feed-forward layers, while the text transformer architecture is based on BERT Devlin et al. [2018]. Q-Former employs a cross-attention mechanism where the query vectors selectively attend to different aspects of the molecule features, allowing the model to capture critical details necessary for understanding and generating textual descriptions of molecular properties.
Specifically, we incorporate causal masks into the queries, ensuring that they only interact with preceding queries. This ensures the sequence of queries maintains a causal dependency, aligning with the requirements of LLMs operating on text sequence. The Causal Q-Former is illustrated in Figure 4. We employ the Causal Q-Former to generate causal queries $\mathbf{Z}=\{\boldsymbol{z}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times d}$ containing high-level molecular and textual information, where $M$ represents the number of queries and $d$ denotes the dimension of queries. Next, we introduce three tailored objectives MTC, MTM, and MTG for the pretraining of the Causal Q-Former.
Figure 4: Illustration of our proposed Causal Q-Former. The Causal Q-Former provides causal queries for subsequent blocks.
Molecule-Text Contrastive Learning (MTC) aims to align molecule and text features by maximizing their mutual information. This is achieved by maximizing the molecule-text similarity of positive pairs against that of negative pairs. We utilize the last query $\boldsymbol{z}_{M}$ of the query sequence $\{\boldsymbol{z}_{i}\}_{i=1}^{M}$ as the query representation, since the output query sequence is causal and the last query contains global information from the queries. For text representation, we use the output embedding of the [CLS] token, denoted as $\boldsymbol{y}$. The contrastive learning loss is expressed as follows:
$\mathcal{L}_{\text{MTC}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp((\boldsymbol{z}_{M}^{i})^{T}\boldsymbol{y}^{i}/\tau)}{\sum_{j=1}^{B}\exp((\boldsymbol{z}_{M}^{i})^{T}\boldsymbol{y}^{j}/\tau)}\\ -\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp((\boldsymbol{y}^{i})^{T}\boldsymbol{z}_{M}^{i}/\tau)}{\sum_{j=1}^{B}\exp((\boldsymbol{y}^{i})^{T}\boldsymbol{z}_{M}^{j}/\tau)},$ (4)
where $B$ denotes the batch size, and $\tau$ represents the temperature parameter. Here, $\boldsymbol{z}_{M}^{i}$ and $\boldsymbol{y}^{i}$ refer to the $i$-th query representation and text representation in a batch, respectively.
Molecule-Text Matching (MTM) focuses on learning fine-grained alignment between molecule and text features. As queries $\{\boldsymbol{z}_{i}\}_{i=1}^{M}$ capture both molecular and textual information through cross-attention and self-attention layers respectively, we utilize the last query $\boldsymbol{z}_{M}$ as input to a binary classifier. This classifier predicts whether a given molecule-text pair is matched or unmatched. The corresponding loss function is formulated as follows:
$\mathcal{L}_{\text{MTM}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\phi(\boldsymbol{z}_{M}\mid\mathbf{X}^{i},\boldsymbol{t}^{i}))}{\sum_{j=1}^{B}\exp(\phi(\boldsymbol{z}_{M}\mid\mathbf{X}^{i},\boldsymbol{t}^{j}))+\sum_{j=1}^{B}\exp(\phi(\boldsymbol{z}_{M}\mid\mathbf{X}^{j},\boldsymbol{t}^{i}))}.$
where $\phi$ represents a binary classifier, and $\mathbf{X}^{i}$ and $\boldsymbol{t}^{i}$ denote the $i$-th input molecule features and input text in a batch, respectively.
Molecule-grounded Text Generation (MTG) focuses on generating textual descriptions given a molecule input. In this task, causal masks for queries are not applied since only textual output is required. However, causal masks are applied for text, allowing each text token to attend to its preceding text tokens and all queries, but not subsequent tokens. The Language Modeling (LM) loss function is applied to model the generation of text $\boldsymbol{t}^{i}$ conditioned on the molecule input $\mathbf{X}^{i}$, formulated as:
$\mathcal{L}_{\text{MTG}}=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{L}\log p\left(t^{i}_{j}\mid t^{i}_{1},\cdots,t^{i}_{j-1},\mathbf{X}^{i}\right),$ (6)
where $t^{i}_{j}$ represents the $j$-th token in the text sequence $\boldsymbol{t}^{i}$. Here, $\mathbf{X}^{i}$ and $\boldsymbol{t}^{i}$ denote the $i$-th input molecule features and generated text in a batch, respectively.
The total loss for training the Causal Q-Former encompasses the three aforementioned objectives:
$\mathcal{L}_{\text{Q-Former}}=\mathcal{L}_{\text{MTC}}+\mathcal{L}_{\text{MTM}}+\mathcal{L}_{\text{MTG}}.$ (7)
Appendix B Details of Datasets
Dataset Type Tasks Description URL License
BBBP Classification Molecular Property Prediction Predicts blood-brain barrier penetration ability. BBBP URL CC-BY 4.0
Tox21 Classification Molecular Property Prediction Toxicity prediction using the Tox21 initiative data. Tox21 URL Public Domain
ToxCast Classification Molecular Property Prediction Broad toxicity prediction with various biological assays. ToxCast URL Public Domain
Sider Classification Molecular Property Prediction Predicts drug side effects. Sider URL CC-BY 4.0
ClinTox Classification Molecular Property Prediction Clinical trial toxicity prediction. ClinTox URL Public Domain
MUV Classification Molecular Property Prediction Virtual screening for unbiased validation. MUV URL CC-BY 4.0
HIV Classification Molecular Property Prediction Predicts anti-HIV activity of molecules. HIV URL Public Domain
BACE Classification Molecular Property Prediction Predicts inhibitors of the BACE-1 enzyme. BACE URL Public Domain
QM9 Regression Molecular Property Prediction Predicts various molecular properties such as atomization energy, dipole moment, etc. QM9 URL CC-BY 4.0
PubChem Captioning, Retrieval, Generation Molecule Captioning, Molecule-Text Retrieval, Caption-guided Molecule Generation Generates descriptions, retrieves text / molecules based on input molecules / text, and guides molecule generation from captions. PubChem URL Public Domain
ChEBI-20 Captioning Molecule Captioning Generates detailed descriptions of molecular structures. ChEBI-20 URL CC-BY 4.0
PCdes Retrieval Molecule-Text Retrieval Used for evaluating accuracy in molecule-text retrieval tasks. PCdes URL CC-BY 4.0
MoMu Retrieval Molecule-Text Retrieval Dataset for molecule-text interaction and retrieval evaluation. MoMu URL CC-BY 4.0
USPTO Generation Reagent Prediction, Forward Reaction Prediction, Retrosynthesis Provides data for predicting reagents, forward reaction outcomes, and retrosynthetic pathways. USPTO URL CC-BY 4.0
Table 5: Summary of datasets, their types, tasks, descriptions, URLs, and licenses used for evaluating UniMoT.
This section provides detailed information about the datasets used in evaluating the performance of UniMoT across various tasks. The datasets are utilized for molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis task. Each dataset serves a unique purpose in assessing different capabilities of the model. We provide a comprehensive overview of datasets, including their types, associated tasks, descriptions, URLs for access, and licensing information.
We present the details of the Molecular Property Prediction Datasets below:
  • BBBP Wu et al. [2018]: The Blood-Brain Barrier Penetration dataset predicts the ability of molecules to penetrate the blood-brain barrier.
  • Tox21 Wu et al. [2018]: This dataset is part of the Toxicology in the 21st Century initiative, used for toxicity prediction.
  • ToxCast Wu et al. [2018]: Another toxicity prediction dataset with a broader range of biological assays.
  • Sider Wu et al. [2018]: Side Effect Resource database, used for predicting drug side effects.
  • ClinTox Wu et al. [2018]: Clinical Toxicity dataset for predicting clinical trial toxicity outcomes.
  • MUV Wu et al. [2018]: Maximum Unbiased Validation dataset for virtual screening.
  • HIV Wu et al. [2018]: Human Immunodeficiency Virus dataset for predicting anti-HIV activities.
  • BACE Wu et al. [2018]: Beta-Secretase 1 dataset for predicting inhibitors of the BACE-1 enzyme, relevant for Alzheimer’s research.
  • QM9 Fang et al. [2023]: The quantum mechanics properties dataset, where the objective is to predict key quantum mechanics properties of a given molecule, such as HUMO, LUMO, and the HUMO-LUMO gap.
We present the details of the Molecule Captioning Datasets below:
  • PubChem Kim et al. [2023]: A large dataset of chemical molecules used for generating textual descriptions of molecular structures.
  • ChEBI-20 Edwards et al. [2022]: A subset of the Chemical Entities of Biological Interest database, provides structured and detailed descriptions of molecules.
We present the details of the Molecule-Text Retrieval Datasets below:
  • PubChem Kim et al. [2023]: Used for both molecule-to-text (M2T) and text-to-molecule (T2M) retrieval tasks.
  • PCdes Zeng et al. [2022]: Another dataset for evaluating M2T and T2M retrieval accuracy.
  • MoMu Su et al. [2022]: Dataset specifically designed for molecule-text interactions and retrieval tasks.
We present the details of the Molecule Generation Datasets below:
  • Mol-Instructions Fang et al. [2023]: This benchmark includes tasks such as caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. It is used to evaluate the model’s ability to generate molecular structures based on textual descriptions and other related tasks.
  • PubChem Kim et al. [2023]: Used for caption-guided molecule generation, generating molecular structures based on textual descriptions.
  • USPTO Fang et al. [2023]: Used for reagent prediction, forward reaction prediction, and retrosynthesis, providing data for predicting reagents, reaction outcomes, and retrosynthetic pathways.
We summarize the datasets used for evaluating UniMoT in Table 5. It encompasses various types of datasets, including those for classification, regression, captioning, retrieval, and generation tasks. Each dataset is described in terms of its type, tasks it supports, a brief description of its content, its URL for access, and the license under which it is distributed. The licenses vary, with some datasets being in the public domain and others under CC-BY 4.0 license.
Appendix C Details of Training
Stage-1: Causal Q-Former Pretraining.
During Stage-1, we only connect the molecule encoder and the Causal Q-Former, leaving out other blocks. We leverage the pretrained molecule encoder from MoleculeSTM Liu et al. [2023a], which has undergone extensive contrastive learning with molecule-text pairs. We utilize the PubChem Kim et al. [2023] dataset for pretraining, keeping the molecule encoder frozen while updating only the Causal Q-Former. Both queries and text serve as input to the Causal Q-Former, while only queries serve as input in subsequent stages. Inspired by BLIP-2 Li et al. [2023], we employ three tailored objectives – Molecule-Text Contrastive Learning (MTC), Molecule-Text Matching (MTM), and Molecule-grounded Text Generation (MTG) – for the pretraining of the Causal Q-Former, as detailed in Appendix A.
The dimension of molecule features is set to 300. We use 16 queries, each with a dimension of 768. The size of $\mathbf{Z}$ ($16\times 768$) is much smaller than the size of molecule features $\mathbf{X}$ (e.g., $150\times 300$). The Q-former is pretrained for 50 epochs. We adopt the AdamW optimizer with a weight decay of 0.05, and a cosine decay learning rate scheduler, with a minimal learning rate of 1e-5. The batch size is set to 64. The computational overhead for this pretraining is 20 GPU hours on 4 NVIDIA A100 GPUs.
Stage-2: Molecule Tokenizer Pretraining.
We connect the Causal Q-Former with the subsequent blocks and train the molecule tokenizer using the objective defined in Equation (2). We utilize SMILES strings Weininger [1988] to represent molecules, and employ the pretrained ChemFormer Irwin et al. [2022] as the generative model. Specifically, we leverage the SMILES encoder and SMILES decoder components provided by ChemFormer. We utilize PubChem Kim et al. [2023] and ChEBI-20 Edwards et al. [2022] datasets, keeping the molecule encoder, SMILES encoder, and SMILES decoder frozen, while updating the Causal Q-Former, codebook, and adapter. Once optimized, the molecule tokenizer remains unchanged throughout the subsequent stages.
The molecule codebook size is set to $K=2048$, and the dimension of codebook embedding is 768. The tokenizer is pretrained for 50 epochs. We adopt the AdamW optimizer with a weight decay of 0.05, and a cosine decay learning rate scheduler, with a minimal learning rate of 1e-5. The batch size is set to 64. The computational overhead for this pretraining is 40 GPU hours on 4 NVIDIA A100 GPUs.
Stage-3: Unified Molecule-Text Pretraining.
We connect the molecule tokenizer with the LLM and employ the LM objective defined in Equation (3) to pretrain the LLM. We utilize Llama Touvron et al. [2023] as the default LLM. To construct the unified molecule-text vocabulary, we merge 2048 molecule codes with the original text vocabulary. Pretraining the LLM involves molecule-to-text autoregression and text-to-molecule autoregression, aimed at enhancing UniMoT’s multi-modal comprehension and generation capabilities. We utilize datasets PubChem Kim et al. [2023] and ChEBI-20 Edwards et al. [2022] for this purpose. To enhance efficiency, we train the LLM using LoRA Hu et al. [2021].
The multi-modal LLM is pretrained for 10 epochs. We adopt the AdamW optimizer with a weight decay of 0.05, and a cosine decay learning rate scheduler, with a minimal learning rate of 1e-5. The batch size is set to 32. The computational overhead for this pretraining is 50 GPU hours on 4 NVIDIA A100 GPUs. To reduce CUDA memory usage, we integrate LoRA with the parameters set to $r=8$, $\alpha=32$, and dropout = 0.1. This integration is applied to the k_proj, v_proj, q_proj, and o_proj modules.
Stage-4: Task-Specific Instruction Tuning.
We perform instruction tuning to align UniMoT with human instructions through supervised fine-tuning on seven tasks: molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. For the molecular property prediction task, we utilize the quantum mechanics properties dataset Fang et al. [2023] for regression prediction and the MoleculeNet Wu et al. [2018] datasets for property classification. For the molecule captioning and molecule-text retrieval tasks, we employ datasets PubChem Kim et al. [2023], PCdes Zeng et al. [2022], and MoMu Su et al. [2022]. For the molecule generation tasks, we utilize the Mol-Instructions Fang et al. [2023] benchmark to conduct instruction tuning. We fine-tune UniMoT for 10 epochs on each task using the same optimizer, learning rate scheduler, and LoRA configurations as in Stage-3 pretraining. Instruction samples for comprehension and generation tasks are shown in Table 6.
Task Instruction
Molecular Property Prediction (Regression) Instruction: Could you give me the LUMO energy value of this molecule?
(Optional: The SMILES sequence is: SMILES)
Output: 0.0576.
Molecular Property Prediction (Classification) Instruction: Evaluate whether the given molecule is able to enter the blood-brain barrier.
(Optional: The SMILES sequence is: SMILES)
Output: Yes.
Molecule Captioning Instruction: Could you give me a brief overview of this molecule?
(Optional: The SMILES sequence is: SMILES)
Output: The molecule is an indole phytoalexin that …
Molecule-Text Retrieval Instruction: Retrieve relevant text for the given molecule.
(Optional: The SMILES sequence is: SMILES)
Output: The molecule is associated with …
Caption-Guided Molecule Generation Instruction: Create a molecule with the structure as described: The molecule is a primary arylamine that …
Output: SMILES of the molecule.
Reagent Prediction Instruction: Please provide possible reagents based on the following chemical reaction.
<REACTANT A> <REACTANT B> ... >> <PRODUCTs>
Output: SMILES of the reagents.
Forward Reaction Prediction Instruction: With the provided reactants and reagents, propose potential products:
<REACTANT A> <REACTANT B> ... <REAGENT A> <REAGENT B> ...
Output: SMILES of the products.
Retrosynthesis Instruction: Please suggest potential reactants and reagents used in the synthesis of the products: <PRODUCTs>
Output: SMILES of the reactants and reagents.
Table 6: Instruction samples for comprehension and generation tasks: molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis.
We have summarized the detailed training hyperparameters of UniMoT in Table 7.
Configuration Q-Former Pretraining Tokenizer Pretraining LLM Pretraining
Molecule Encoder MoleculeSTM MoleculeSTM MoleculeSTM
SMILES Encoder - ChemFormer ChemFormer
SMILES Decoder - ChemFormer ChemFormer
LLM Base - - Llama-2-7B
Epoch 50 50 10
Optimizer AdamW AdamW AdamW
Codebook Size 2048 2048 2048
Number of Queries 16 16 16
Query Emb. Dim. 768 768 768
Molecule Emb. Dim. 300 300 300
Batch Size 64 64 32
Minimal LR 1e-5 1e-5 1e-5
LR Scheduler Cosine Cosine Cosine
Warm-up Steps 1000 1000 1000
Weight Decay 0.05 0.05 0.05
LoRA Config - - $r=8$, $\alpha=32$, $\text{dropout}=0.1$
Precision bfloat16 bfloat16 bfloat16
GPU Usage 4 NVIDIA A100s 4 NVIDIA A100s 4 NVIDIA A100s
Training Time 20 GPU hours 40 GPU hours 50 GPU hours
Table 7: The detailed training hyperparameters of UniMoT.
Appendix D Details and More Results of Experiments
Molecular Property Prediction Task.
Property prediction aims to anticipate a molecule’s intrinsic physical and chemical properties based on its structural or sequential characteristics. In the regression task, we conduct experiments on the quantum mechanics properties dataset QM9 Fang et al. [2023], where the objective is to predict key quantum mechanics properties of a given molecule, such as HUMO, LUMO, and the HUMO-LUMO gap. We compare UniMoT against several baselines, including Alpaca Taori et al. [2023], Baize Xu et al. [2023], Llama-2-7B Touvron et al. [2023], Vicuna-13B Chiang et al. [2023], Mol-Instructions Fang et al. [2023], and InstructMol Cao et al. [2023]. Mean Absolute Error (MAE) serves as our evaluation metric. The performance of the regression task on the QM9 dataset is presented in Table 8. Compared to previous single-modal instruction-tuned LLMs and molecular LLMs, UniMoT exhibits further improvement on the regression task, showcasing its fundamental comprehension abilities in molecular contexts.
[…]
Alpaca (Llama-7B) - - - 322.109
Baize (Llama-7B) - - - 261.343
Llama-2-7B 0.7367 0.8641 0.5152 0.7510
Vicuna-13B 0.7135 3.6807 1.5407 1.9783
Mol-Instructions (Llama-7B) 0.0210 0.0210 0.0203 0.0210
InstructMol (Vicuna-7B) 0.0048 0.0050 0.0061 0.0050
UniMoT (Llama-2-7B) 0.0042 0.0047 0.0055 0.0049
Table 8: Mean Absolute Error (MAE) of molecular property prediction task (regression) on the QM9 Fang et al. [2023] dataset. Bold indicates the best performance and underline indicates the second best performance. $\Delta\epsilon$ is the HOMO-LUMO energy gap.
Molecule Captioning Task.
The molecule captioning task involves generating a comprehensive description of a molecule. For this task, we compare UniMoT with several baselines: MolT5 Edwards et al. [2022], MoMu Su et al. [2022], InstructMol Cao et al. [2023], MolCA Liu et al. [2023b], and 3D-MoLM Li et al. [2024]. We adopt BLEU Papineni et al. [2002], ROUGE Lin [2004], and METEOR Banerjee and Lavie [2005] as the evaluation metrics. The performance of UniMoT in the molecule captioning task on the ChEBI-20 Edwards et al. [2022] dataset is presented in Table 9. Some concrete examples of molecule captioning task are presented in Table 10. From the results, it is evident that UniMoT consistently outperforms the baselines by a significant margin. These results underscore the effectiveness of the molecule tokenizer in providing molecule tokens with high-level molecular and textual information, thus enhancing molecule comprehension.
[…]
T5-Small 50.1 41.5 60.2 44.6 54.5 53.2
T5-Base 51.1 42.3 60.7 45.1 55.0 53.9
T5-Large 55.8 46.7 63.0 47.8 56.9 58.6
MolT5-Small (T5-Small) 51.9 43.6 62.0 46.9 56.3 55.1
MolT5-Base (T5-Base) 54.0 45.7 63.4 48.5 57.8 56.9
MolT5-Large (T5-Large) 59.4 50.8 65.4 51.0 59.4 61.4
MoMu-Small (T5-Small) 53.2 44.5 - - 56.4 55.7
MoMu-Base (T5-Base) 54.9 46.2 - - 57.5 57.6
MoMu-Large (T5-Large) 59.9 51.5 - - 59.3 59.7
InstructMol (Vicuna-7B) 47.5 37.1 56.6 39.4 50.2 50.9
MolCA (OPT-125M) 61.6 52.9 67.4 53.3 61.5 63.9
MolCA (OPT-1.3B) 63.9 55.5 69.7 55.8 63.6 66.9
UniMoT (Llama-2-7B) 66.4 58.3 72.2 58.4 66.4 70.3
Table 9: Performance (%) of molecule captioning task on the ChEBI-20 Edwards et al. [2022] dataset. Bold indicates the best performance and underline indicates the second best performance.
Molecule Generated Molecule Caption Ground Truth
[Uncaptioned image] The molecule is an optically active form of phenylalaninate having D-configuration. It is a conjugate base of a D-phenylalanine. It is an enantiomer of a L-phenylalaninate. The molecule is the D-enantiomer of phenylalaninate. It is a conjugate base of a D-phenylalanine. It is an enantiomer of a L-phenylalaninate.
[Uncaptioned image] The molecule is an ammonium ion that is the conjugate acid of 2-phenylpropylamine arising from protonation of the primary amino function; major species at pH 7.3. It has a role as a human metabolite, an Escherichia coli metabolite and a mouse metabolite. It is a conjugate acid of a 2-phenylpropylamine. The molecule is the cation obtained by protonation of the amino group of 2-phenylethylamine. It has a role as a human metabolite and an Escherichia coli metabolite. It is a conjugate acid of a 2-phenylethylamine.
[Uncaptioned image] The molecule is an enamide obtained by the carboxy group of trans-cinnamic acid with the secondary amino group of (2S,5R)-1,2,5-trimethylpiperazine. It has a role as an Aspergillus metabolite. It is an alkaloid, a N-acylpiperazine, an enamide and a tertiary carboxamide. It derives from a trans-cinnamic acid. The molecule is an enamide obtained by formal condensation of the carboxy group of trans-cinnamic acid with the secondary amino group of (2R,5R)-1,2,5-trimethylpiperazine. It has a role as an Aspergillus metabolite. It is a N-acylpiperazine, a N-alkylpiperazine, an alkaloid, an enamide and a tertiary carboxamide. It derives from a trans-cinnamic acid.
[Uncaptioned image] The molecule is an (omega-1)-hydroxy fatty acid ascaroside obtained by formal condensation of the alcoholic hydroxy group of (10R)-10-hydroxylauric acid with ascarylopyranose (the alpha anomer). It is a metabolite of the nematode Caenorhabditis elegans. It has a role as a Caenorhabditis elegans metabolite. It is a monocarboxylic acid and an (omega-1)-hydroxy fatty acid ascaroside. It derives from an (11R)-11-hydroxylauric acid. It is a conjugate acid of an ascr18(1-). The molecule is an (omega-1)-hydroxy fatty acid ascaroside obtained by formal condensation of the alcoholic hydroxy group of (10R)-10-hydroxyundecanoic acid with ascarylopyranose (the alpha anomer). It is a metabolite of the nematode Caenorhabditis elegans. It is a monocarboxylic acid and an (omega-1)-hydroxy fatty acid ascaroside. It derives from a (10R)-10-hydroxyundecanoic acid. It is a conjugate acid of an ascrblue18(1-).
[Uncaptioned image] The molecule is a 2-oxo monocarboxylic acid that is pyruvic acid in which one of the methyl hydrogens is substituted by a 4-vinylcyclohex-2-en-1-yl group. It has a role as a plant metabolite. It derives from a pyruvic acid. It is a conjugate acid of a 4-[(1E)-4-vinylcyclohex-2-en-1-yl]pyruvate. The molecule is a 2-oxo monocarboxylic acid that is pyruvic acid in which one of the methyl hydrogens has been replaced by a methylenecyclopropyl group. It has a role as a rat metabolite and a xenobiotic metabolite. It is a 2-oxo monocarboxylic acid, a member of cyclopropanes and an olefinic compound. It derives from a pyruvic acid.
Table 10: Examples of molecule captioning task on the ChEBI-20 dataset. We highlight in blue the text that accurately describes the molecule structures in the generated caption, ensuring alignment with the ground truth.
Molecule-Text Retrieval Task.
Model Retrieval in batch Retrieval in test set
M2T (%) T2M (%) M2T (%) T2M (%)
[…]
Sci-BERT 85.3 98.7 84.2 98.4 41.7 87.3 40.2 86.8
KV-PLM 86.1 98.6 85.2 98.5 42.8 88.5 41.7 87.8
MoMu (Sci-BERT) 87.6 99.2 86.4 99.4 47.3 90.8 48.1 89.9
MoMu (KV-PLM) 88.2 99.4 87.3 99.4 48.5 91.6 49.5 90.7
MoleculeSTM 90.5 99.6 88.6 99.5 52.7 92.9 53.2 92.5
MolCA (OPT-1.3B) 92.6 99.8 91.3 99.5 67.9 94.4 68.6 93.3
3D-MoLM (Llama-2-7B) 93.5 100.0 92.9 99.6 69.1 95.9 70.1 94.9
UniMoT (Llama-2-7B) 93.6 100.0 92.7 99.4 69.5 96.3 69.8 94.4
Table 11: Performance (%) of molecule-text retrieval task on the PubChem Kim et al. [2023] dataset. Bold indicates the best performance and underline indicates the second best performance.
Model Retrieval in batch Retrieval in test set
M2T (%) T2M (%) M2T (%) T2M (%)
Sci-BERT 62.6 61.8 60.7 60.8
KV-PLM 77.9 65.0 75.9 64.3
MoMu (Sci-BERT) 80.6 77.0 79.1 75.5
MoMu (KV-PLM) 81.1 80.2 80.2 79.0
MoleculeSTM 86.2 83.9 84.6 85.1
MolCA (OPT-1.3B) 91.4 88.4 90.5 87.6
3D-MoLM (Llama-2-7B) 92.3 89.6 91.2 88.5
UniMoT (Llama-2-7B) 92.6 89.4 91.6 88.3
(a) Accuracy (%) of molecule-text retrieval task on the PCdes Zeng et al. [2022] dataset.
Model Retrieval in batch Retrieval in test set
M2T (%) T2M (%) M2T (%) T2M (%)
Sci-BERT 1.4 1.6 0.3 0.3
KV-PLM 1.5 1.3 0.5 0.3
MoMu (Sci-BERT) 45.7 40.0 43.3 43.4
MoMu (KV-PLM) 46.2 38.5 43.7 43.5
MoleculeSTM 81.8 81.9 75.8 74.5
MolCA (OPT-1.3B) 83.7 84.3 88.6 87.3
3D-MoLM (Llama-2-7B) 84.9 85.4 89.9 88.7
UniMoT (Llama-2-7B) 85.4 85.6 90.3 89.0
(b) Accuracy (%) of molecule-text retrieval task on the MoMu Su et al. [2022] dataset.
Table 12: Accuracy (%) of molecule-text retrieval task on the PCdes Zeng et al. [2022] and MoMu Su et al. [2022] datasets. Bold indicates the best performance and underline indicates the second best performance. We report the performance of retrieval using a batch of 64 random samples and the entire test set.
The molecule-text retrieval task involves using a molecule to retrieve text (M2T) and using text to retrieve a molecule (T2M). We compare UniMoT with several baselines: Sci-BERT Beltagy et al. [2019], KV-PLM Zeng et al. [2022], MoMu Su et al. [2022], MoleculeSTM Liu et al. [2023a], MolCA Liu et al. [2023b], and 3D-MoLM Li et al. [2024]. We report the performance of retrieval using a batch of 64 random samples and the entire test set, evaluated with the metrics of Accuracy and Recall@20. We use the checkpoint from Stage-1 of pretraining. UniMoT is evaluated on the datasets of PubChem Kim et al. [2023], PCdes Zeng et al. [2022], and MoMu Su et al. [2022]. Performance on the PubChem dataset is shown in Table 11, while performance on the PCdes Zeng et al. [2022] and MoMu Su et al. [2022] datasets is shown in Table 12. UniMoT demonstrates superior performance over the baselines on molecule-text retrieval, particularly in molecule-to-text retrieval. This demonstrates that UniMoT has learned fine-grained alignment between molecules and text, and it can understand molecule-text interactions through the introduction of the Causal Q-Former.
Molecule Generation Tasks.
Molecule generation tasks include caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. Caption-guided molecule generation involves creating molecular structures from textual descriptions, leveraging NLP and cheminformatics to interpret and translate descriptions into chemical structures. Reagent prediction focuses on identifying suitable reagents for given reactants and desired products, optimizing synthetic routes. Forward reaction prediction forecasts probable products from specific reactants and reagents, using knowledge of chemical reactivity. Retrosynthesis deconstructs target molecules into simpler starting materials. The results of forward reaction prediction and retrosynthesis deconstructs are presented in Table 13.
In molecule generation tasks, evaluating the quality of generated molecules involves several metrics that measure different aspects of similarity and validity. Exact Match checks if the generated molecule is identical to the target molecule, offering a stringent criterion for precise replication but potentially overlooking chemically similar variants. The BLEU score Papineni et al. [2002], adapted from machine translation, measures the overlap of n-grams (short sequences of atoms or bonds) between generated and target molecules, thus assessing partial similarities. Levenshtein Distance Levenshtein and others [1966] evaluates the minimum number of edits needed to transform the generated molecule into the target, providing insight into structural changes required. RDKit Landrum and others [2006], MACCS Durant et al. [2002], and Morgan Morgan [1965] Fingerprint Similarities compare the generated and target molecules based on various molecular fingerprinting methods, which capture different aspects of molecular structure and properties. The Validity Kusner et al. [2017] metric assesses the proportion of chemically valid molecules generated, ensuring that the output consists of plausible chemical structures. Together, these metrics offer a comprehensive evaluation framework, balancing exact matches with structural and chemical validity.
[…]
Forward Reaction Prediction
Llama 0.000 0.020 42.002 0.001 0.002 0.001 0.039
Vicuna 0.000 0.057 41.690 0.007 0.016 0.006 0.059
Mol-Instructions 0.045 0.654 27.262 0.313 0.509 0.262 1.000
InstructMol 0.536 0.967 10.851 0.776 0.878 0.741 1.000
UniMoT 0.611 0.980 8.297 0.836 0.911 0.807 1.000
Retrosynthesis
Llama 0.000 0.036 46.844 0.018 0.029 0.017 0.010
Vicuna 0.000 0.057 46.877 0.025 0.030 0.021 0.017
Mol-Instructions 0.009 0.705 31.227 0.283 0.487 0.230 1.000
InstructMol 0.407 0.941 13.967 0.753 0.852 0.714 1.000
UniMoT 0.478 0.974 11.634 0.810 0.909 0.771 1.000
Table 13: Performance of molecule generation tasks on the Mol-Instructions Fang et al. [2023] benchmark, including forward reaction prediction and retrosynthesis. Bold indicates the best performance, and underline indicates the second best performance.
Appendix E Additional Ablation Studies
LLM Architecture and Adaptation.
We conducted a comparison of molecule captioning performance across various LLM architectures and adaptation strategies, as illustrated in Table 14. Our experiments show that UniMoT performs well across multiple LLM architectures, including Galactica Taylor et al. [2022] and Mistral Jiang et al. [2023] series, demonstrating its robustness and generalizability. The experiments also indicate that scaling up the LLM to 13B or adopting a full fine-tuning (FFT) strategy yields only marginal improvements in performance compared to using Llama-2-7B with LoRA. While larger models and FFT strategy might offer slight gains in performance, they come at a significant cost in terms of efficiency.
Architecture Adaptation BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
Galactica-125M LoRA 28.7 21.5 34.2 21.1 30.3 31.0
Galactica-1.3B LoRA 30.2 22.8 36.0 22.4 32.2 33.2
Mistral-7B LoRA 32.0 24.2 38.0 24.0 34.1 35.2
Llama-2-7B LoRA 31.3 23.8 37.5 23.7 33.6 34.8
Llama-2-7B FFT 32.0 24.6 38.3 24.3 34.7 35.6
Llama-2-13B LoRA 31.8 24.3 38.0 24.1 34.4 35.3
Table 14: Ablation study on the LLM architecture and adaptation strategy.
Models with Comparable Sizes.
We conducted a comprehensive performance comparison between UniMoT and MolCA Liu et al. [2023b] using models of comparable sizes, as detailed in Table 15. The results show that UniMoT consistently outperforms MolCA across various LLM architectures, including Galactica-125M, Galactica-1.3B, and LLaMA-2-7B. This consistent performance highlights the effectiveness of UniMoT in handling molecule-to-text tasks, further validating the superiority of tokenizer-based architecture over adapter-based architecture. The tokenizer-based architecture can achieve better molecule-text alignment through autoregressive molecule-to-text and text-to-molecule pretraining compared to other architectures.
Model BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
MolCA (Galactica-125M) 25.9 17.5 34.4 16.6 23.9 28.5
MolCA (Galactica-1.3B) 28.6 21.3 36.2 21.4 29.7 32.6
MolCA (Llama-2-7B) 28.2 21.0 33.5 20.9 30.0 30.8
UniMoT (Galactica-125M) 28.7 21.5 34.2 21.1 30.3 31.0
UniMoT (Galactica-1.3B) 30.2 22.8 36.0 22.4 32.2 33.2
UniMoT (Llama-2-7B) 31.3 23.8 37.5 23.7 33.6 34.8
Table 15: Performance of UniMoT and MolCA using comparable model sizes on the molecule captioning task using the PubChem dataset.
Query Size.
We also conducted an ablation study to evaluate the performance of UniMoT with different query sizes, as presented in Table 16. The results indicate that increasing the query size leads to improved performance, with the best performance achieves at a query size of 32. However, this larger query size also demands significantly more training time and memory. Therefore, for a more efficient balance between performance and resource consumption, we opt to use a query size of 16, which still offers strong performance while being more computationally feasible.
Architecture Query Size BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
Llama-2-7B 4 25.1 18.3 30.2 18.5 26.1 27.3
Llama-2-7B 8 29.5 21.3 34.5 21.8 30.9 31.5
Llama-2-7B 16 31.3 23.8 37.5 23.7 33.6 34.8
Llama-2-7B 32 32.2 24.9 38.2 24.4 34.7 35.7
Table 16: Performance of UniMoT with different query sizes on the molecule captioning task using the PubChem dataset.
Appendix F Experimental Results with Additional Baselines
We aim to enhance the molecule comprehension and generation experiments in the main text by including additional baselines such as EdgePred Hu et al. [2019], GraphCL You et al. [2020], Mole-BERT Xia et al. [2022], MoMu Su et al. [2022], GIT-Mol Liu et al. [2024b], and MolCA Liu et al. [2023b]. These baselines are presented in Tables 17, 18, 19, and 20.
[…]
KV-PLM 70.50 72.12 55.03 59.83 89.17 54.63 65.40 78.50
EdgePred 67.30 76.00 64.10 60.40 64.10 74.10 76.30 79.90
AttrMask 67.79 75.00 63.57 58.05 75.44 73.76 75.44 80.28
InfoGraph 64.84 76.24 62.68 59.15 76.51 72.97 70.20 77.64
MolCLR 67.79 75.55 64.58 58.66 84.22 72.76 75.88 71.14
GraphMVP 68.11 77.06 65.11 60.64 84.46 74.38 77.74 80.48
GraphCL 69.70 73.90 62.40 60.50 76.00 69.80 78.50 75.40
Mole-BERT 71.90 76.80 64.30 62.80 78.90 78.60 78.20 80.80
MoMu-S 70.50 75.60 63.40 60.50 79.90 70.50 75.90 76.70
MoMu-K 70.10 75.60 63.00 60.40 77.40 71.10 76.20 77.10
MoleculeSTM 69.98 76.91 65.05 60.96 92.53 73.40 76.93 80.77
GIT-Mol 73.90 75.90 66.80 63.40 88.30 - - 81.08
InstructMol (Vicuna-7B) 70.00 74.67 64.29 57.80 91.48 74.62 68.90 82.30
MolCA (OPT-1.3B) 70.00 77.20 64.50 63.00 89.50 - - 79.80
UniMoT (Llama-2-7B) 71.37 76.43 65.78 59.79 92.89 75.97 78.49 83.69
Table 17: ROC-AUC (%) of molecular property prediction task (classification) on the MoleculeNet Wu et al. [2018] datasets. Bold indicates the best performance and underline indicates the second best performance.
[…]
MolT5-Small (T5-Small) 22.5 15.2 30.4 13.5 20.3 24.0
MolT5-Base (T5-Base) 24.5 16.6 32.2 14.0 21.4 26.1
MolT5-Large (T5-Large) 25.9 17.3 34.1 16.4 23.4 28.0
MoMu-Small (T5-Small) 22.9 16.0 31.0 13.7 20.8 24.4
MoMu-Base (T5-Base) 24.7 16.8 32.5 14.6 22.1 27.2
MoMu-Large (T5-Large) 26.3 18.0 34.8 16.9 24.8 28.7
InstructMol (Vicuna-7B) 18.9 11.7 27.3 11.8 17.8 21.3
MolCA (OPT-125M) 25.9 17.5 34.4 16.6 23.9 28.5
MolCA (OPT-1.3B) 28.6 21.3 36.2 21.4 29.7 32.6
3D-MoLM (Llama-2-7B) 30.3 22.5 36.8 22.3 31.2 33.1
UniMoT (Llama-2-7B) 31.3 23.8 37.5 23.7 33.6 34.8
Table 18: Performance (%) of molecule captioning task on the PubChem Kim et al. [2023] dataset. Bold indicates the best performance and underline indicates the second best performance.
[…]
T5-Small 50.1 41.5 60.2 44.6 54.5 53.2
T5-Base 51.1 42.3 60.7 45.1 55.0 53.9
T5-Large 55.8 46.7 63.0 47.8 56.9 58.6
MolT5-Small (T5-Small) 51.9 43.6 62.0 46.9 56.3 55.1
MolT5-Base (T5-Base) 54.0 45.7 63.4 48.5 57.8 56.9
MolT5-Large (T5-Large) 59.4 50.8 65.4 51.0 59.4 61.4
MoMu-Small (T5-Small) 53.2 44.5 - - 56.4 55.7
MoMu-Base (T5-Base) 54.9 46.2 - - 57.5 57.6
MoMu-Large (T5-Large) 59.9 51.5 - - 59.3 59.7
InstructMol (Vicuna-7B) 47.5 37.1 56.6 39.4 50.2 50.9
MolCA (OPT-125M) 61.6 52.9 67.4 53.3 61.5 63.9
MolCA (OPT-1.3B) 63.9 55.5 69.7 55.8 63.6 66.9
UniMoT (Llama-2-7B) 66.4 58.3 72.2 58.4 66.4 70.3
Table 19: Performance (%) of molecule captioning task on the ChEBI-20 Edwards et al. [2022] dataset. Bold indicates the best performance and underline indicates the second best performance.
[…]
Caption-guided Molecule Generation
Llama 0.000 0.003 59.864 0.005 0.000 0.000 0.003
Vicuna 0.000 0.006 60.356 0.006 0.001 0.000 0.001
Mol-Instructions 0.002 0.345 41.367 0.231 0.412 0.147 1.000
MolT5 0.112 0.546 38.276 0.400 0.538 0.295 0.773
UniMoT 0.237 0.698 27.782 0.543 0.651 0.411 1.000
Reagent Prediction
Llama 0.000 0.003 28.040 0.037 0.001 0.001 0.001
Vicuna 0.000 0.010 27.948 0.038 0.002 0.001 0.007
Mol-Instructions 0.044 0.224 23.167 0.237 0.364 0.213 1.000
InstructMol 0.129 0.610 19.664 0.444 0.539 0.400 1.000
UniMoT 0.167 0.728 14.588 0.549 0.621 0.507 1.000
Forward Reaction Prediction
Llama 0.000 0.020 42.002 0.001 0.002 0.001 0.039
Vicuna 0.000 0.057 41.690 0.007 0.016 0.006 0.059
Mol-Instructions 0.045 0.654 27.262 0.313 0.509 0.262 1.000
InstructMol 0.536 0.967 10.851 0.776 0.878 0.741 1.000
UniMoT 0.611 0.980 8.297 0.836 0.911 0.807 1.000
Retrosynthesis
Llama 0.000 0.036 46.844 0.018 0.029 0.017 0.010
Vicuna 0.000 0.057 46.877 0.025 0.030 0.021 0.017
Mol-Instructions 0.009 0.705 31.227 0.283 0.487 0.230 1.000
InstructMol 0.407 0.941 13.967 0.753 0.852 0.714 1.000
UniMoT 0.478 0.974 11.634 0.810 0.909 0.771 1.000
Table 20: Performance of molecule generation tasks on the Mol-Instructions Fang et al. [2023] benchmark, including caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. Bold indicates the best performance, and underline indicates the second best performance.
Appendix G Broader Impacts
The development of UniMoT, a unified model for molecule and text modalities, has significant potential to positively impact various fields. UniMoT can streamline the drug discovery process by enabling efficient molecule generation and optimization based on textual descriptions. In material science, it can aid in discovering new materials with desirable properties. Additionally, UniMoT can enhance research collaboration between chemists, biologists, and data scientists by integrating molecular and textual data, leading to comprehensive research insights and innovative solutions.
This paper does not pose any ethical concerns. The study does not involve human subjects and follows proper procedures for dataset releases. There are no potentially harmful insights, methodologies, or applications. Additionally, there are no conflicts of interest or sponsorship concerns. Discrimination, bias, and fairness issues are not applicable. Privacy and security matters have been appropriately addressed, legal compliance has been maintained, and research integrity has been upheld.